OpenStudy (anonymous):

Solve using the given root:2x3 + 9x2 + 7x - 6 = 0; -2

OpenStudy (cwrw238):

first divide (x + 2) into the polynomial to get a quadratic equation

OpenStudy (cwrw238):

use long or synthetic division.

OpenStudy (anonymous):

these are the anwers to choose from a{-2,0.5, 3}, b{-2,0.5,-3} , c {-2,-0.5,-3} ,d {-2,1,0}

OpenStudy (cwrw238):

you would learn more if you did the division are you familiar with long division of polynomials?

OpenStudy (anonymous):

some what

OpenStudy (cwrw238):

i'll start this one off for you * 2x^2 + 5x ------------------- * x + 2 ) 2x^3 + 9x^2 + 7x - 6 (2x^3 / x = 2x^2, 2x^2 * 2 = 4x^2 ** ) 2x^3 + 4x^2 ** ----------- subtract: 5x^2 + 7x 5x^2 + 10x (5x^2 / x = 5x, 5x * 2 = 10x) -----------

OpenStudy (cwrw238):

do u see how i got 4x^2 and 10 x the quotient is above the top line